OpenStudy (anonymous):

Friends, is this right? Simplify the quantities in Exercise using m(z) = z^2. 1) m(z+h)-m(z-h) (z^2+h)-(z^2-h) =0

OpenStudy (anonymous):

m(z+h)-m(z-h) \[(z^2+h)-(z^2-h) = 0\]

OpenStudy (mathstudent55):

\(m(z) = z^2\) \(m(z + h) - m(z - h)\) \(= (z + h)^2 - (z - h)^2\) \(= z^2 + 2hz + h^2 - z^2 + 2hz - h^2\) \(= 4hz\)

OpenStudy (anonymous):

@mathstudent55 How did we get the square root outside the brackets? I appreciate your help.

OpenStudy (mathstudent55):

In this problem, there are no square roots. there is an exponent of 2.

OpenStudy (anonymous):

oh sorry, exponent :)))

OpenStudy (mathstudent55):

This is how I interpreted this problem. The line m(z) = z^2 means function m of variable z is defined as z^2.

OpenStudy (mathstudent55):

Then you are asked to do m(z + h) - m(z - h) This means evaluate function m at z + h, and subtract from it function m evaluated at z - h.

OpenStudy (mathstudent55):

Since function m(z) equals z^2, then function m(z) evaluated at z + h means to replace z with z + h. You get: m(z) = z^2 m(z + h) = (z + h)^2 Similarly, you must evaluate function m(z) at z - h, so replace z with z - h. You get: m(z) = z^2 m(z - h) = (z - h)^2 This means that m(z + h) - m(z - h) = (z + h)^2 - (z - h)^2 Now you need to square the binomials z + h and z - h, and do the subtraction. (z + h)^2 - (z - h)^2 = = z^2 + 2hz + h^2 - (z^2 - 2hz + h^2) = z^2 + 2hz + h^2 - z^2 + 2hz - h^2 = 4hz

zepdrix (zepdrix):

Sometimes color helps if you're finding this too confusing c:\[\large\rm m(\color{orangered}{z})=(\color{orangered}{z})^2\]So then,\[\large\rm m(\color{orangered}{z+h})=(\color{orangered}{z+h})^2\]

zepdrix (zepdrix):

Just trying to further explain why the square is on the outside c:

OpenStudy (anonymous):

do you have an m(h) ?

OpenStudy (victoriasushchik):

m(z+h) = (z+h)² = z² + 2zh + h² m(z-h) = (z-h)² = z² - 2zh + h² m(z+h) - m(z-h) = z² + 2zh + h² - (z² - 2zh + h²) = 4zh